OpenStudy (anonymous):

Is my answer correct? Please help me!!! xoxo (7-2√ 5) (5 + 4√ 5)

OpenStudy (anonymous):

\[35 + 20 sq \root 5\]

OpenStudy (anonymous):

\[35 + 20 sq \root 5\]

OpenStudy (anonymous):

Not quite... you missed some of the parts of the multiplication.

OpenStudy (anonymous):

[-5+18\sqrt{5}\]

OpenStudy (anonymous):

I got -5 + 18 square root 5....

OpenStudy (anonymous):

\[-5+18\sqrt{5}\]

OpenStudy (anonymous):

You have to distribute when you multiply... (7-2√ 5) (5 + 4√ 5) = 7*5 + 7*4√ 5 - 5*2√ 5 - (2√ 5) (4√ 5)

OpenStudy (anonymous):

I didn't break it down enough

OpenStudy (anonymous):

thank you guys!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! :) <3

OpenStudy (anonymous):

good work!... glad you solved it... even better that you figured it out on your own :)

OpenStudy (anonymous):

thank you jakeypoo!!!! :)@JakeV8